OpenStudy (anonymous):

Solve r^2 - 4r = 21 by completing the square.

OpenStudy (anonymous):

\[r^2-4r=21\\ r^2-4r+4=21+4\\ (r-2)^2=25\\ r-2=\pm\sqrt{25}\\ ~~~~~~~~~\vdots\]